OpenStudy (anonymous):

what is the square root of 1792

TheSmartOne (thesmartone):

Welcome to Open Study!!

OpenStudy (confluxepic):

@chrstnr It is better to explain it. I know there is nothing to explain in this question but it may be helpful to confused students.

TheSmartOne (thesmartone):

\[\sqrt{1792}=\sqrt{256\times7}=\sqrt{256}\times \sqrt{7}=16\sqrt{7}\]

TheSmartOne (thesmartone):

\[16\sqrt{7}=42.332020977033449448025852058228\]

TheSmartOne (thesmartone):

It goes on foreever... but if you need it in decimal you can keep it aas 42.33

TheSmartOne (thesmartone):

@tjacqueline6